Eva is risk averse. Currently she has $50,000 to invest. She faces the following

choice: she can invest in the stock of a dotcom company, or she can invest in IBM stock. If she

invests in the dot com company, then with probability 0.5 she will lose $30,000, but with

probability 0.5 she will gain $50,000. If she invests in IBM stock, then with probability 0.5 she

will lose only $10,000, but with probability 0.5 she will gain only $30,000. Can you tell w hich

investment she will prefer to make?


1
Expert's answer
2022-01-04T11:43:19-0500

Expected return from investment in dotcom company:

"E(R)=(0.5\times-30000)+(0.5\times50000)=\$10,000"


Expected return from investment in IBM:

"E(R)=(0.5\times-10000)+(0.5\times30000)=\$10,000"


The expected return from both stocks is same ($10,000). But the possibility of losing money is higher in case of dotcom company. So, Eva should prefer IBM stock for investment. It is because this investment is suited with her risk averse nature.